OpenStudy (anonymous):

Solve 6x2 = 18x. x = 0, x = −3 x = 0, x = 3 x = 6, x = −3 x = 6, x = 3

OpenStudy (aroub):

x = 0, x = 3

OpenStudy (anonymous):

how did u got that aroub?

jimthompson5910 (jim_thompson5910):

6x^2=18x 6x^2-18x=0 6x(x-3)=0 6x=0 or x-3=0 x=0 or x=3

OpenStudy (aroub):

6x^2-18x=0 6x(x-3) so x=0 .x=3

OpenStudy (anonymous):

Thanks!

OpenStudy (aroub):

no prob! aand thank you! :)